# VILA2: VILA Augmented VILA

### v1

@@ Abstract
Equal contribution.
Visual language models (VLMs) have rapidly progressed, driven by the success of large language models (LLMs). While model architectures and training infrastructures advance rapidly, data curation remains under-explored. When data quantity and quality become a bottleneck, existing work either directly crawls more raw data from the Internet that does not have a guarantee of data quality or distills from black-box commercial models (e.g., GPT-4V [1] / Gemini [2]) causing the performance upper bounded by that model. In this work, we introduce a novel approach that includes a self-augment step and a specialist-augment step to iteratively improve data quality and model performance. In the self-augment step, a VLM recaptions its own pretraining data to enhance data quality, and then retrains from scratch using this refined dataset to improve model performance. This process can iterate for several rounds. Once self-augmentation saturates, we employ several specialist VLMs finetuned from the self-augmented VLM with domain-specific expertise, to further infuse specialist knowledge into the generalist VLM through task-oriented recaptioning and retraining. With the combined self-augmented and specialist-augmented training, we introduce VILA2 (VILA-augmented-VILA), a VLM family that consistently improves the accuracy on a wide range of tasks over prior art, and achieves new state-of-the-art results on MMMU leaderboard [3] among open-sourced models.
@@ 1 Introduction
The success of large language models (LLMs) [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] has offered the cornerstone for cross-modality tasks. Through the alignment of visual encoders to LLMs, visual language models have enabled myriad appealing capabilities to visual tasks, such as instruction following, zero-shot generalization, few-shot in-context learning (ICL), and enhanced world knowledge [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. The field has progressed rapidly in the past two years, yielding effective alignment training recipes [18, 23, 25] and model architectures [16, 17, 18, 19, 20].
@@
Contrary to the fast-evolving training enhancement, the underlying human-generated datasets and tasks remain simple [26, 27, 28, 29]. Given the costly nature of VLM training, most methods are confined with coarse-quality large-scale captioning image-text pairs (pretraining), followed by fine-grained small-scale supervised finetuning (SFT). Enhancement of image-text pairs with millions and billions of instances can inevitably impose a huge amount of human effort, and thus not realistic. Recent methods have observed rewarding distillation possibilities from proprietary commercial models like GPT-4V [1] and Gemini [2]. However, the performance is upper bounded by these models. In the meantime, studies remain very sparse on how to better utilize VLMs to correct human error and remedy dataset task simplicity for enhanced training.
@@
In this work, we aim to answer “whether it is possible that the VLM itself can remedy dataset deficiency and enhance its training.” We delve deep into the potential of using VLM itself to refine and augment the pretraining data and performance in an iterative way. Our new training regime is summarized in Figure 1, which consists of two main steps: a self-augment step and a specialist-augment step. We start with the self-augment loop (Figure 1 (a)) that leverages VLMs to enhance the quality of pretraining data. We demonstrate that synthetic data, combined with the original data, can collaboratively generate stronger models in a bootstrapped loop manner. Intuitively and as we observed, the loops offer performance boosts for free, but suffers diminishing returns after 3 rounds. To facilitate additional learning, we reformulate a more challenging task-specific loop (Figure 1 (b)). In this loop, a specialist with a focus on new knowledge or tasks, such as a specially-aware expert, is finetuned from the self-augmented VLM using a limited amount of additional SFT data. The specialist can then recaption a massive amount of pretraining data. Finally, the self-augmented VLM can be retrained on the specialist-recaptioned pretraining data to further boost the performance. We demonstrate the benefits of this loop using three different specialist VLMs with expertise in spatial awareness, OCR, and grounding.
@@
Our novel VLM augmentation training regime progressively improves data quality, i.e., covering enhanced visual semantics and reducing hallucinations, as demonstrated in Figure 1 and more in our experimental section. This offers a direct performance boost to VLMs. We introduce a new family of VILA2 models, as in VILA-augmented-VILA. VILA2 outperforms state-of-the-art methods across main benchmarks, all enhanced without bells and whistles via self-bootstrapped training. Now constituting a new state of the art on MMMU benchmark [3] among all open-sourced models, we hope that the insights and release of VILA2’s recipe, data, and code can assist with our community for better understanding and usage of synthetic data to train stronger VLMs.
@@
Figure 1: The schematic diagram to train VILA2, short for VILA-augmented VILA. We re-formulate visual language model (VLM) training with model in the loop to remedy training data defficacy. We start with validating design options in constructing a self-augmenting loop (Section. 2.1) to improve on caption quality of the default training task. After the saturation of this process, we challenge the VLM to generate data conforming to extra SFT-enabled tasks to further VLM learning (Section. 2.2). Our new design insights yield off-the-shelf performance boosts to VLMs (Section. 3).
@@ 2 Methodology
In this paper, we focus on auto-regressive VLMs where image tokens are projected into the textual space and concatenated with text tokens, in line with [16, 2, 25]. This approach is chosen because of its flexibility when handling multi-modal inputs. We follow the widely adapted three-stage training paradigm, i.e., align-pretrain-SFT, to ablate our studies. We start to self-augment VLM training by constructing a bootstrapped loop leveraging VLM’s general captioning capability. After the bootstrapping saturates, we then introduce specialist augmenting exploiting VLM’s skills picked up during SFT across additional visual tasks as specialist feedback to its pretraining stage. We next elaborate on these steps in detail.
@@ 2.1 Self-augmenting via General Knowledge Enhancement
MMC4 [26] COYO [28] COYO-$\text{VILA}_{1}$ COYO-$\text{VILA}_{2}$ COYO-$\text{VILA}_{3}$ COYO-$\text{VILA}_{4}$
Avg #Words 17.1 ± 25.0 11.9 ± 9.0 101.2 ± 43.0 117.1 ± 49.1 126.77 ± 50.10 125.9 ± 51.2
@@
Table 1: The average number of words and $\text{VQA}^{\text{v2}}$ evaluation comparison between the original dataset and the re-captioned dataset. Best performance is bolded and second best is underlined. During self-augmentation, the caption lengths increases significantly, thus offering more details and information. The length and VQA score growth plateau after three rounds of self-augmenting.
Existing VLM training largely relies on data from the internet, where the texts are usually brief and short, see Table. 1 where the average number of words is less than 20 for MMC4 [26] and COYO [28]. In addition to briefness, human annotations can also fall short in explaining to LLMs the versatile semantics an image presents. As another example, Figure. 4 indicates that an original COYO caption only describes the person riding on the street while lacking a description of clothing and surroundings. Previous studies have either assigned humans to write dense captions or relied on commercial propriety APIs for image descriptions. The first option can be labor-intensive and costly, while the second may inherit model biases, upper bound the model performance, and also potentially introduce copyright issues. Such detailed information, however, can be highly beneficial for VLM training, as we show later.
Rather than distilling proprietary models or relying on manual laboring, we aim to use VILA to generate better captions for VILA’s pre-training, exploiting the power of the already-intelligent VILA within intermediate training stages to conduct laborious relabelling effort.
Beginning with the original dataset, we first train the initial version of VILA, referred to as $\text{VILA}_{0}$ in subsequent experiments. Next, we use $\text{VILA}_{0}$ to re-caption VILA’s pre-training datasets. With appropriate prompt choice and conversation template,  $\text{VILA}_{0}$ is able to generate long and detailed captions. Then the augmented datasets, consisting of real images from the internet and synthetic texts from  $\text{VILA}_{0}$, are then re-used to train the next round of VILA, named $\text{VILA}_{1}$. This self-augmenting process can be repeated several rounds before convergence, as shown in Figure 1 (a).
Figure 2: LLM judgement for captions from $\text{VILA}_{i}$, $i$ indicating the self-augmenting round. Evaluations are based on 5,000 sampled data from Coyo [28]. Both GPT-4V [1] and Gemini-1.0-pro [2] prefer for VILA2 augmented texts and captions from later rounds got higher score.
@@ 2.1.1 Prompts and Template Design
We also augment prompt styles in the self-augmentation loop to diversify caption styles. This approach leads to immediate performance improvements. The key idea is to advance the knowledge embedded in new captions and to avoid repetitive generations.
To validate design choices, we conducted an in-depth study on prompt choices as follows, where <img> indicates the location where image features will be inserted, and discuss our findings.
@@
Brief and Short Re-captioning is NOT Helpful. We begin with a straightforward prompt asking VLMs to briefly describe the image. Despite these brief recaptions being significantly longer than the original texts (90 vs. 17 words), there is no notable improvement in VLM benchmarks, as shown in Table. 1. In fact, metrics even deteriorate in benchmarks such as Science-Image and MMMU. This decline is due to the lack of details and the potential hallucinations introduced during recaption.
Next, we re-design the prompt to encourage VLMs to provide a more detailed description of visual narrative elements in images. We also referenced the prompt template from ShareGPT-4V [30] to ensure the descriptions are accurate and precise. Our experiments demonstrate that using three different long prompts improves the quality of recaptioning and boosts performance in benchmarks, with a clear ranking among them, detailed in Table. 1. Therefore, we leverage a mixture of these prompts by randomly selecting from versions v1 to v3.
Keeping Origin Human Text is Important.
We compare different conversation templates in Table. 1, where the first only uses real human data and concatenated adapts both human and synthetic description. We conduct experiments to train with both choices and find that using self-augmented data leads to improvement on major benchmarks such as LLaVA-Bench, Science, TextVQA, and MME. However, we notice there is still a decline in several metrics, prompting us to concatenate both the origin and re-captioned texts in order to best preserve information. This template consistently improves all VLM metrics (Table. 1).
With the proper prompt choice and template design, self-augmenting can continuously the caption quality. Shown in Table. 1, the synthetic captions become much longer and more informative than real human texts (reflecting on VQAv2 score). We further follow LLMJudge [11] to compare caption quality from different rounds. Both Gemini-1.0-Pro and GPT-4V prefer VILA2 generated captions (higher winrate).
@@ 2.2 Surpassing the Limit with Specialist VLM Augmentation
Figure 3: VILA2 specialist VLM augmentation overview. We start by gathering task-specific knowledge to create task-specialist VLMs. These specialist VLMs annotate images with task-oriented prompts. The generated question-answering pairs are then utilized to re-train the next iteration of VILA2.
We observe that the improvement from self-augmentation starts to saturate after a few rounds of bootstrapping, with all free lunches having been squeezed (Table. 3). We hypothesize that this shortcoming stems from the monotonic task of general descriptive captioning, which is also heavily influenced by language modeling priors.
To further facilitate the focus of visual features and enable effective annotation, we propose incorporating extra task-specific knowledge into the generalist VLM to create several specialist VLMs. Each specialist model is finetuned with data that requires a deeper understanding of image components and semantics, e.g., spatial relations, localization, and OCR.
It might seem as if we require a separate specialist model for each task whereas we show that quick finetuning of VLM during the SFT stage towards downstream tasks already suffices for this role. A bootstrapped loop can then transfer such specialist knowledge from small-scale SFT data onto a large number of pretraining images. In this manner, we jointly exploit the synergy of abundant image amounts in pretraining and specialist knowledge acquired during SFT.
@@ 2.2.1 Acquiring Specialized Knowledge
In this paper, we focus on three challenging tasks that require direct visual understanding: spatial relations understanding, Grounded narration, and OCR, as summarized in Figure 3. We next elaborate on the process of specialist construction as follows:
Spatial Specialist.
@@
To explicitly challenge the model to acquire additional spatial awareness, we curated SpatialRelationQA, a dataset containing 1 million conversations about spatial relations within scenes. Our dataset is built on LV3D, a comprehensive collection of both indoor and outdoor 3D datasets from Cube-LLM [31] that is designed to enhance the understanding of 3D spatial relations requiring both perceptual and grounding skills.
We formulated a three-step process to create the QA pairs. First, for each cleaned 3D scene, we iterated through all 3D bounding boxes and randomly sampled from five object-object relations (closest, in front of, behind, left, right), with six object-camera relations (close, far, closest, farthest, left, right). Next, we checked if any remaining bounding boxes matched these sampled relations. Finally when a match was found, we randomly selected question templates to construct the QA pairs, incorporating instances, their projected 2D bounding boxes, and relations. A sample question can be constructed as: “Where is the chair closest to the table [$x_{\text{left}}$,$y_{\text{top}}$,$x_{\text{right}}$,$y_{\text{bottom}}$] in the image?”, with answer being the target bounding box.
Grounding Specialist.
@@
To also enhance knowledge towards grounding awareness we exploited grounded narration, a highly visual-centric task that requires VLMs to generate detailed captions to accurately locate major visual elements using 2D bounding boxes, as shown in Figure 3. This approach provides dense supervision and allows us to verify if VLMs hallucinate. To develop the VILA2 grounding specialist, we used image-grounded caption pairs from the 20M GRIT dataset [32]. We first filtered out bounding boxes covering more than 70% of the image area, as many images in GRIT are book or album covers not closely related to the captions. Next, we removed images containing more than three instances of the same category to reduce complexity and decrease noise in generation orders. This process yielded 4M higher-quality instances for grounding specialist training, which we further split into two subsets: Grounding-Short 3M and Grounding-Long 838K for a two-stage finetuning.
OCR Specialist.
@@
Finally we focus on OCR strength enrichment. We utilize a diverse set of images featuring textual content, such as tables, charts, and documents, to develop an OCR specialist. Each image is annotated with QA pairs that focus on text recognition (e.g., Q: What is the title of the book?), comprehension (e.g., Q: Which bar has the largest value?), and reasoning (e.g., Q: What is the main idea of the quote from Albert Camus?). Dataset details are provided in appendix A.2.
@@
The three specialist are then applied to the final augmentation stage. To this end we use a new set of task-oriented prompts to activate the specialists’ knowledge and improve their instruction-following ability by narrowing focus. Specifically, we prompt the spatial relations understanding specialist with evenly sampled templates of "<image> Elaborate on the visual and narrative elements of the image in detail, with a focus on spatial relations." and "<image> Can you explain the content of the image and their spatial relations in detail?" during the specialist augmentation stage. Similarly, the grounding specialist generated captions with bounding boxes for the major visual focus, and the OCR specialist identified most textual content in the images. The responses from these different specialist VLMs are then curated as QA pairs and appended to the original captions of COYO images for the next pertaining iteration of VILA2. More details can be found in appendix A.1.

### v2

# VILA2: VLM Augmented VLM with Self-Improvement

@@ Abstract
While visual language model (VLM) architectures and training infrastructures advance rapidly, data curation remains under-explored where quantity and quality become a bottleneck. Existing work either crawls extra Internet data with a loose guarantee of quality or distills from black-box proprietary models (e.g., GPT-4V / Gemini) that are API frequency and performance bounded. This work enables a VLM to improve itself via data enhancement, exploiting its generative nature. We introduce a simple yet effective VLM augmentation scheme that includes a self-augment step and a specialist-augment step to iteratively improve data quality and hence, model performance. In the self-augment step, the instruction-finetuned VLM recaptions its pretraining caption datasets and then retrains from scratch leveraging refined data. Without any expensive human-in-the-loop annotation, we observe improvements in data quality and downstream accuracy boosts with three self-augmentation rounds – a viable free lunch to the current VLM training recipe. When self-augmentation saturates, we augment the caption diversity by leveraging specialty skills picked up from instruction finetuning. We finetune VLM specialists from the self-augmented VLM with domain-specific experts, including spatial, grounding, and OCR, to fuse task-aware synthetic data into the pretraining stage. Data quality improvements and hallucination reductions are cross-checked by VLM (GPT-4V, Gemini) and human judges. Combining self-augmentation and specialist-augmented training, VILA2 consistently improves the accuracy on a wide range of benchmarks over the prior art, producing a reusable pretraining dataset that is 300x more cost-efficient than human labeling.
@@ 1 Introduction
The success of large language models (LLMs) (Raffel et al., 2020; Dai et al., 2019; Brown et al., 2020; OpenAI, 2023b; Touvron et al., 2023a; b; Taori et al., 2023; Chiang et al., 2023; Karamcheti et al., 2021; Chowdhery et al., 2022; yi, 2023; Bai et al., 2023a) has offered the cornerstone for cross-modality tasks. Through the alignment of visual encoders to LLMs, visual language models have enabled myriad appealing capabilities to visual tasks, such as instruction following, zero-shot generalization, few-shot in-context learning (ICL), and enhanced world knowledge (Liu et al., 2023c; Alayrac et al., 2022; Driess et al., 2023; Chen et al., 2023c; Li et al., 2023b; fuy, 2023; Bai et al., 2023b; OpenAI, 2023a; Zhu et al., 2023a; Lin et al., 2024). The field has progressed rapidly in the past two years, yielding effective alignment training recipes (Driess et al., 2023; OpenAI, 2023a; Lin et al., 2024) and model architectures (Liu et al., 2023c; Alayrac et al., 2022; Driess et al., 2023; Chen et al., 2023c; Li et al., 2023b).
@@
Contrary to the fast-evolving training enhancement, the underlying human-generated datasets and tasks remain simple (Zhu et al., 2023b; Schuhmann et al., 2022; Byeon et al., 2022; Sharma et al., 2018). Given the costly nature of VLM training, most methods are confined with coarse-quality large-scale captioning image-text pairs (pretraining), followed by fine-grained small-scale supervised finetuning (SFT). Enhancement of image-text pairs with millions and billions of instances can inevitably impose a huge amount of human effort, and thus not realistic. Recent methods have observed rewarding distillation possibilities from proprietary commercial models like GPT-4V (OpenAI, 2023c) and Gemini (Gemini Team & other authors, 2023). However, the performance is upper bounded by these models. In the meantime, studies remain very sparse on how to better utilize VLMs to correct human error and remedy dataset task simplicity for enhanced training.
@@
In this work, we aim to answer “whether it is possible that the VLM itself can remedy dataset deficiency and enhance its training.” We delve deep into the potential of using VLM itself to refine and augment pretraining data and performance iteratively. Our new training regime, summarized in Figure 1, consists of two main steps: a self-augment step and a specialist-augment step. We start with the self-augment loop (Figure 1 (a)) that leverages VLMs to enhance the quality of pretraining data. We demonstrate that synthetic data, combined with the original data, can collaboratively generate stronger models in a bootstrapped loop manner. Intuitively and as we observed, the loops offer performance boosts for free, but suffers diminishing returns after 3 rounds. To facilitate further learning, we reformulate a more challenging task-specific loop (Figure 1 (b)). In this loop, specialists with a focus on new knowledge or tasks, such as a spatial-aware expert, OCR expert or grounding expert, are finetuned from the self-augmented VLM using a limited amount of additional SFT data. The specialists can then recaption a massive amount of pretraining data. Finally, the self-augmented VLM can be retrained on the specialist-recaptioned pretraining data to further boost the performance.
@@
The insights yield a novel VLM augmentation training regime progressively improves data quality, by transferring knowledge from the higher-quality but small-scale SFT stage back to the larger-scale but coarse pretraining phase. This improvements cover enhanced visual semantics (Figure 1) and reducing hallucinations (Table 3.43.4). We also observed consistent agreements on data quality improvements when cross checking the data by VLM models (GPT-4V, Gemini) and humans (Ph.D. students). This offers a direct performance boost to VLMs. We introduce a new family of VILA2 models, as in VLM-augmented-VLM. VILA2 outperforms state-of-the-art methods across main benchmarks, all enhanced without bells and whistles via self-bootstrapped training. We hope that the insights and release of VILA2’s recipe, data, and code can assist with our community for better understanding and usage of synthetic data to train stronger VLMs.
@@
Figure 1: The schematic diagram to train VILA2, short for VILA-augmented VILA. We re-formulate visual language model (VLM) training with “model in the loop” to remedy training data defficacy. We start with validating design options in constructing a self-augmenting loop (Section. 2.1) to improve on caption quality of the default training task. After the saturation of this process, we challenge the VLM to generate data conforming to extra SFT-enabled tasks to further VLM learning (Section. 2.2). Our new design insights yield off-the-shelf performance boosts to VLMs (Section. 3).
@@ 2 Methodology
In this paper, we focus on auto-regressive VLMs where image tokens are projected into the textual space and concatenated with text tokens, in line with (Liu et al., 2023c; Gemini Team & other authors, 2023; Lin et al., 2024). This approach is chosen because of its flexibility when handling multi-modal inputs. We follow the widely adapted three-stage training paradigm, i.e., align-pretrain-SFT, to ablate our studies. We start to self-augment VLM training by constructing a bootstrapped loop leveraging VLM’s general captioning capability. After the bootstrapping saturates, we then introduce specialist augmenting exploiting VLM’s skills picked up during SFT across additional visual tasks as specialist feedback to its pretraining stage. We next elaborate on these steps in detail.
@@ 2.1 Self-augmenting via General Knowledge Enhancement
Existing VLM training largely relies on data from the Internet, where the texts are usually brief and short, see Table 1 where the average number of words is less than 20 for MMC4 (Zhu et al., 2023b) and COYO (Byeon et al., 2022). In addition to brevity, human annotations can also fall short in explaining to LLMs the versatile semantics an image presents. As another example, Figure. 4 indicates that an original COYO caption that only describes a person riding on the street, omitting details about clothing and surroundings. Previous studies have either assigned humans to write dense captions or by using commercial propriety APIs for detailed descriptions. The first option can be labor-intensive and costly, while the second risks model biases, limiting model performance, and potentially raising copyright concerns.
Figure 2: LLM judgement for captions from $\text{VILA}_{i}$, $i$ indicating the self-augmenting round. Evaluations are based on 5,000 sampled data from Coyo Byeon et al. (2022). Both GPT-4V OpenAI (2023c) and Gemini-1.0-pro Gemini Team & other authors (2023) prefer for VILA2 augmented texts and captions from later rounds got higher scores.
MMC4 COYO COYO-$\text{VILA}_{1}$ COYO-$\text{VILA}_{2}$ COYO-$\text{VILA}_{3}$ COYO-$\text{VILA}_{4}$
Avg #Words 17.1 $\pm$ 25.0 11.9 $\pm$ 9.0 101.2 $\pm$ 43.0 117.1 $\pm$ 49.1 126.77 $\pm$ 50.10 125.9 $\pm$ 51.2
@@
Table 1: The average number of words and $\text{VQA}^{\text{v2}}$ evaluation comparison between the original dataset and the re-captioned dataset. Best performance is bolded and second best is underlined. During self-augmentation, the caption lengths increases significantly, thus offering more details.
Rather than distilling proprietary models or relying on manual laboring, we aim to use VILA to generate better captions for VILA’s pretraining. This approach exploits the power of the already-intelligent VILA within intermediate training stages to conduct laborious relabelling efforts. We begin with the original dataset to train the initial version of VILA, referred to as $\text{VILA}_{0}$ in subsequent experiments. Next, we use $\text{VILA}_{0}$ to re-caption VILA’s pretraining datasets. With appropriate prompt choice and conversation template,  $\text{VILA}_{0}$ is able to generate long and detailed captions. Then the augmented datasets, consisting of real images from the internet and synthetic texts from  $\text{VILA}_{0}$, are used to train the next round of VILA, named $\text{VILA}_{1}$. This self-augmenting process can be repeated several rounds before convergence, leading to detailed descriptions (higher $\text{VQA}^{\text{v2}}$ score in Table 1) and improved text quality (LLM Judge in Figure. 2).
@@ 2.1.1 Prompts and Template Design
The choice of prompt is particularly important for immediate performance improvements. To validate prompt design choices, we conducted an in-depth study on prompt choices as follows and discuss our findings, where <img> indicates the location where image features will be inserted,
@@
Brief and Short Re-captioning is NOT Helpful. We begin with a straightforward prompt asking VLMs to briefly describe the image. Despite these brief recaptions being significantly longer than the original texts (90 vs. 17 words), there is no notable improvement in VLM benchmarks, as shown in Table 2. In fact, metrics even deteriorate in benchmarks such as Science-Image and MMMU-Test. This decline may stem from a lack of details during recaption.
Next, we redesign the prompt to encourage VLMs to provide a more detailed description of visual narrative elements in images. We also referenced the prompt template from ShareGPT-4V (Chen et al., 2023b) to ensure the descriptions are accurate and precise. Our experiments demonstrate that using three different long prompts improves the quality of recaptioning and boosts performance in benchmarks, detailed in Table 1. Therefore, we leverage a mixture of these prompts by randomly selecting from versions v1 to v3.
Keeping Original Human Text is Important. We compare different conversation templates in Table 1. The first template uses only real human data, while the “concatenated" approach adapts both human and synthetic descriptions. Our experiments reveal that using self-augmented data improves performance on major benchmarks like LLaVA-Bench, Science-Image, TextVQA. However, we noticed a decline in several metrics. This prompted us to concatenate both the original and re-captioned texts to best preserve information, which consistently improves all VLM metrics. (Table 1).
@@ 2.2 Surpassing the Limit with Specialist VLM Augmentation
While self-augmentation provides a simple yet effective way to boost VLM’s performance, we notice that the improvement starts to saturate with all free lunches having been squeezed (Table 3). We hypothesize that this shortcoming stems from the monotonic task of general descriptive captioning, which is also heavily influenced by language modeling priors.
To advance the boundary of self-augmentation, we propose the integration of extra task-specific knowledge into a generalist VLM to create several specialist VLMs. Each specialist model is finetuned with data that demands a deeper understanding of image components and semantics, e.g., spatial relations, localization, and OCR. A bootstrapped loop can then transfer such specialist knowledge from small-scale SFT data onto a large number of pretraining images.
@@ 2.2.1 Acquiring Specialized Knowledge
We focus on three challenging tasks: spatial relations understanding, grounded narration, and OCR (Figure 3), and then elaborate the specialist construction as follows:
@@
Spatial Specialist. To explicitly challenge the model to acquire additional spatial awareness, we curated SpatialRelationQA, a dataset containing 1 million conversations about spatial relations within scenes. Our dataset is built on LV3D, a comprehensive collection of both indoor and outdoor 3D datasets from Cube-LLM (Cho et al., 2024) that is designed to enhance the understanding of 3D spatial relations requiring both perceptual and grounding skills.
We formulated a two-step process to create the QA pairs.
  1. 1.
    For each cleaned 3D scene, we iterated through all 3D bounding boxes and randomly sampled from object-object relations (closest, in front of, behind, left, right) and object-camera relations (close, far, closest, farthest, left, right);
  2. 2.
    Next we checked if any remaining bounding boxes matched these sampled relations. For each matched results, we randomly selected question templates to construct the QA pairs, incorporating instances, their projected 2D bounding boxes, and relations.
Figure 3: VILA2 specialist VLM augmentation overview. We gather task-specific knowledge to create task-specialist VLMs. These specialist VLMs annotate images with task-oriented prompts and generate question-answering pairs to re-train the next iteration of VILA2.
A sample question can be constructed as: “Where is the chair closest to the table [$x_{\text{left}}$,$y_{\text{top}}$,$x_{\text{right}}$,$y_{\text{bottom}}$] in the image?”, with answer being the target bounding box. During recaptioning, we guide the model to answer with more spatial information by prompting the specialist VLM with evenly sampled templates during the recaptioning phase, e.g., “<image> Can you explain the content of the image and their spatial relations in detail?”.
@@
Grounding Specialist. To enhance knowledge of grounding awareness, we exploited grounded narration – a highly visual-centric task that requires VLMs to generate detailed captions to accurately locate major visual elements using 2D bounding boxes, as shown in Figure 3. This approach provides dense supervision and allows us to verify if VLMs hallucinate. To develop the VILA2 grounding specialist, we used image-grounded caption pairs from the 20M GRIT dataset (Peng et al., 2023). We first filtered out bounding boxes covering more than 70% of the image area, as many images in GRIT are book or album covers unrelated to the captions. We then removed images containing more than three instances of the same category to reduce complexity and decrease noise in generation orders. This process yielded 4M high-quality instances for grounding specialist training, which we split into two subsets: Grounding-Short (3M) and Grounding-Long (838K) for two-stage finetuning.
@@
OCR Specialist. For OCR capabilities, we utilize a diverse set of images featuring textual content, such as tables, charts, and documents, to develop an OCR specialist. Each image is annotated with QA pairs that focus on text recognition (e.g., Q: What is the title of the book?), comprehension (e.g., Q: Which bar has the largest value?), and reasoning (e.g., Q: What is the main idea of the quote from Albert Camus?). Dataset details are provided in appendix A.2.
@@
The three specialist are then applied to the final augmentation stage. We use a new set of task-oriented prompts to activate the specialists’ knowledge and improve their instruction-following ability by narrowing focus. Specifically, during the specialist augmentation stage. we prompt with evenly sampled templates of "<image> Elaborate on the visual and narrative elements of the image in detail, with a focus on spatial relations." and "<image> Can you explain the content of the image and their spatial relations in detail?" Similarly, the grounding specialist generates captions with bounding boxes for the major visual focus, while the OCR specialist identifies most textual content in the images. The responses from these different specialist VLMs are appended to the original captions as QA pairs for the next pretraining iteration of VILA2. The full details are attached in appendix A.1.